OpenStudy (anonymous):

Jose is paying for a meal using bills in his wallet. He has five $1 bills and four $5 bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two $5 bills to pay for the meal?

OpenStudy (ganpat):

2/4 = 1/2...

OpenStudy (ganpat):

@mathslover : dude is this right ?? m lil confused..

OpenStudy (ledah):

I'm assuming since this is probability there's a 4 out of 9 chance it will be two 5$ bills.

OpenStudy (ledah):

there's also a 5 out of 9 it will be a $1 bill

OpenStudy (ganpat):

or 2/ 4 * 4/9 = 2/ 9 ?? How abt this ??

OpenStudy (ledah):

well I don't quite think that would be correct. because there are more than 2 $5 bills

OpenStudy (ganpat):

hmmm.. @ganeshie8 ??

OpenStudy (anonymous):

thanks a lo t ganpat

OpenStudy (ganpat):

dude wait.. m not sure on this..

OpenStudy (ledah):

I'm going to assume my attempt to help was ignored.

OpenStudy (anonymous):

i think it's near of the answer 0.6 * 0.6

OpenStudy (anonymous):

it wasnt

OpenStudy (ledah):

This has nothing to do with multiplication....

OpenStudy (anonymous):

divide the problem into two steps. step 1: probability of getting $5 bill out off 9 bills, of which 4 are of its own kind (i.e. $5) probability for that would be 4/9 step 2: probability of getting $5 bill out off 8 bills, of which 3 are of its own kind (i.e. $5) probability for that would be 3/8 now multiply these two figures for answer. it would be 1/6

OpenStudy (ganpat):

@Ledah : ha ha, dont get angry dude.. just trying to make it more clear...

OpenStudy (ledah):

I'm not angry I'm just speaking my mind.

OpenStudy (ganpat):

@vamgadu : 3/ 8 ?? were did this come from ?

OpenStudy (ledah):

Yeah... where did that come from?

OpenStudy (anonymous):

when you have removed one $5 bill successfully, there would only be 8 bills in there. as u have removed a $5 bill out the four, there will only be 3 left. so the probability of getting a $5 bill is 3/8

OpenStudy (ledah):

..... I don't think it means removed.... ermm... I believe that it should be what I said earlier.

OpenStudy (ganpat):

@vamgadu : yes.. it says one at a time..

OpenStudy (ledah):

Confused me.

OpenStudy (anonymous):

okay that's mean multy by 2

OpenStudy (anonymous):

me to do not know what to put as an answer

OpenStudy (anonymous):

for at once time two bills

OpenStudy (ledah):

And now I'm completely... confused. Awesome! -_-

OpenStudy (anonymous):

ahhhhhhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

the question said the bills are taken out one at a time without replacement. so It is presumed that they are removed

OpenStudy (anonymous):

yes

OpenStudy (ledah):

But this is still probability.

OpenStudy (anonymous):

yeah

OpenStudy (ledah):

so you wouldn't remove the numbers just because it says removed.

OpenStudy (anonymous):

you do have to or it will be wrong

OpenStudy (ganpat):

@vamgadu : so that is right.. 4/ 9 * 3/8... 1/6.. i agreee

OpenStudy (ledah):

hmm. very well then.

OpenStudy (anonymous):

so is it % 17

OpenStudy (ganpat):

yes.. 17%

OpenStudy (anonymous):

??

OpenStudy (ganpat):

1/ 6 = 0.1666 = 0.17 = 17 %